[(3x=11)/6]-[(2x+1)/3]=x+5
Please help-I'm stuck.

Answer by THANApHD(104)

U mean
[(3x+11)/6]-[(2x+1)/3]=x+5
[3x+11-2(2x+1)]/6 = x+5
3x+11-4x-2 = 6(x+5)=6x+30
3x-4x-6x = 30+2-11
-7x = 21
x = 21/-7
= -3